# DeePer Insight into Your User: Directed Persona Refinement for Dynamic Persona Modeling

### v1

@@ Abstract
To advance personalized applications such as recommendation systems and user behavior prediction, recent research increasingly adopts large language models (LLMs) for human-readable persona modeling. In dynamic real-world scenarios, effective persona modeling necessitates leveraging streaming behavior data to continually optimize user personas. However, existing methods—whether regenerating personas or incrementally extending them with new behaviors—often fail to achieve sustained improvements in persona quality or future behavior prediction accuracy. To address this, we propose DeePer, a novel approach for dynamic persona modeling that enables continual persona optimization. Specifically, we enhance the model’s direction-search capability through an iterative reinforcement learning framework, allowing it to automatically identify effective update directions and optimize personas using discrepancies between user behaviors and model predictions. Extensive experiments on dynamic persona modeling involving 4,800 users across 10 domains highlight DeePer ’s superior persona optimization capabilities, delivering an impressive 32.2% average reduction in user behavior prediction error over four update rounds—outperforming the best baseline by a remarkable 22.92%. [^1] Resources are available at https://github.com/sheep333c/DEEPER.git.
@@ 1 Introduction
Recent studies increasingly utilize Large Language Models (LLMs) OpenAI (2023); Anthropic (2024); AI@Meta (2024) for human-readable and interpretable persona modeling, advancing personalized applications like recommendation and behavior prediction. However, most research focuses on generating personas from static historical data, which fail to capture dynamic behaviors and evolving preferences in real-world interactive scenarios Wang and Lim (2023); Zhou et al. (2024). This underscores the need for dynamic persona modeling—a pivotal yet underexplored approach that iteratively updates personas using streaming user behavior data to continually enhance their quality.
@@
To address the challenges, we propose DeePer (Directed Persona Refinement), a novel approach for LLM-based dynamic persona modeling. Specifically, we introduce a new paradigm, Persona Refinement (Figure 1), which uses discrepancies between user behaviors and model predictions as stronger update signals to expose deficiencies in personas. To identify effective update directions, we decompose the optimization objective into three direction search goals: Previous Preservation, Current Reflection, and Future Advancement, ensuring stability, adaptability, and task alignment. Based on these goals, we design reward functions for clear and measurable assessments of update directions by comparing predictive errors before and after updates. Finally, we propose an iterative reinforcement learning (RL) framework with two training stages, leveraging self-sampling and DPO fine-tuning to progressively enhance the model’s direction search and persona refinement capabilities, ultimately improving prediction accuracy.
@@ 2.1 Task Formulation
Consider a user $\mathcal{U}$ in domain $\mathcal{X}$. To capture temporal dynamics of user behaviors, we segment user’s online interactions into sequential, time-ordered windows $\mathbf{W}=\{\mathcal{W}_{t}\}_{t=0}^{\mathcal{T}}$. Each window $\mathcal{W}_{t}$ contains $N$ interactions, represented by an item list $\mathbf{I}_{t}=\{j_{t}^{j}\}_{j=1}^{N}$ and the corresponding user behaviors $\mathbf{O}_{t}=\{o_{t}^{j}\}_{j=1}^{N}$. As new data arrives at time $t$, the current window $\mathcal{W}_{t}$ captures interactions from the present period, while $\mathcal{W}_{t-1}$ reflects previous behaviors, and $\mathcal{W}_{t+1}$ outlines future interactions.
@@ 3.4 Iterative Training Framework
DeePer employs an iterative training framework (Figure 2): Iteration 1 fine-tunes the base model to refine initial personas (Model 1), while Iteration 2 further enhances it to refine pre-optimized personas (Model 2). Direct Preference Optimization (DPO)Rafailov et al. (2024) is used to seamlessly integrate rewards into preference pairs, enabling the model to identify better directions through explicit comparisons and supporting scalable iterative fine-tuning.
@@ 4.1 Experiment Setup
We use the frozen, powerful LLM, GPT-4o-miniOpenAI (2023), to generate initial personas and predictions in a zero-shot setting for both training and evaluation, ensuring consistent initial persona quality and unbiased predictions. Additionally, it serves as the backbone for all baselines, offering a robust foundation for comparison.
@@ 5.1 What enables DeePer’s effectiveness
We first evaluate the necessity of direction search by comparing DeePer with frozen models (GPT-4o-mini and Llama-3.1-8B-Instruct (Naive Model)) which refine personas directly. As shown in Figure 5(a), both baselines exhibit significant error increases after refinement, underscoring the critical role of direction search in effective optimization.
@@
Previous Window Prediction ($\varepsilon_{prev|\mathcal{S}_{old/new}}$) - Previous Preservarion
@@
Figure 6: (a) Refinement performance across two RL iterations of DeePer. (b) Comparison of DeePer and fine-tuned IncUpdate (Inc-FT).
Iterative RL enhances persona refinement.
@@ 5.2 Persona Probing
Refined personas from DeePer also enable in-depth, domain-specific exploration. In Book domain, we uncover group-level preferences by clustering final persona embeddings, identifying five user groups characterized by unique high-frequency adjectives (e.g., “romantic” and “practical”) (Figure 7(b)). We also extract domain-specific patterns by organizing high-frequency terms into eight dimensions using GPT-4o (Table 3), highlighting critical factors for modeling Book domain users. These attempts show DeePer ’s potential to support strategic user insights exploration.
@@ 7 Conclusion
In this paper, we introduce DeePer, an effective approach to dynamic persona modeling that leverages iterative reinforcement learning and discrepancy-based refinement to continuously enhance persona quality and predictive accuracy. Comprehensive experiments demonstrate DeePer ’s effectiveness across diverse domains in dynamic user modeling. We hope DeePer marks a significant advancement in personalized applications.
@@ A.3 Persona Update
In this work, the formulation of persona update stage varies by corresponding paradigms. For all baselines, we use GPT-4o-mini as backbone, while for DeePer, we use the fine-tune model via the iterative RL training framework based on Llama3.1-8b-Instruct. For each persona update method, we design corresponding update prompt template as follows.
@@
MovieLens
- 20M Dataset
MovieLens Train/Eval 3000 1000 E
@@
Art Crafts and Sewing Eval - 86 F
@@
Sports and Outdoors Eval - 236 H
@@
Grocery and Gourmet Food Eval - 185 I
@@
Appendix B Dataset Details
B.1 User Details
@@
B.2 Training Data Construction
@@
In Iteration 2, Model 1 is first used to generate outputs for the 10,800 context data points from Iteration 1, completing the first persona update for each user. These results, in turn, are used to construct a second set of 10,800 context data points for the second persona refinement. These are then combined with the 10,800 context data points constructed in the first iteration, resulting in a total of 21,600 context data points for sampling in the second iteration. For each context, 15 candidate personas are again randomly sampled using Model 1, with the same inference parameters as in Iteration 1. The boundaries for positive and negative reward sets are set to 0.5 and 0, with a margin of 0.8. A total of 28,612 new DPO preference pairs are generated. Additionally, 5,000 preference pairs with a margin greater than 0.8 are randomly selected from Iteration 1 to be included in the training set, mitigating the issue of catastrophic forgetting. This results in a total of 33,612 DPO preference pairs, with 10% randomly selected for the validation set, used to train Model 2.
@@
Appendix C Training Details
C.1 Hyperparameter Details
The hyperparameters used in the training process are summarized in Table 12.
C.2 Loss Function Details
@@
The DPO loss optimizes the model by leveraging user preference signals to align persona refinements with higher rewards. The loss is defined as:
@@
Appendix D Dynamic Persona Modeling Task
@@
Grocery and Gourmet Food 1.19 1.20 1.17 1.14 1.13
@@
Art Crafts and Sewing 0.76 0.75 0.74 0.67 0.74
Automative 0.84 0.88 0.86 0.94 0.87
Book 1.00 1.03 0.98 0.97 0.96
@@
Grocery and Gourmet Food 1.19 1.19 1.17 1.18 1.16
@@
Movies and TV 1.12 1.12 1.11 1.10 1.09
Recipe 0.91 0.94 0.87 0.83 0.83
Sports and Outdoors 0.91 0.90 0.89 0.91 0.88
@@
Appendix E What Enables DeePer’s Effectiveness
@@
E.1 Proving Effectiveness of Direction search
@@
E.2 Proving the Effectiveness of Balanced Reward
@@
    E.3 Proving the Effectiveness of Iterative RL Training
    In this analysis, we aim to demonstrate the effectiveness of iterative RL training by comparing the performance of the model after one iteration of refinement versus two iterations. This comparison helps to understand whether additional refinement iterations contribute to improved model performance.
@@ E.3 Proving the Effectiveness of Iterative RL Training
    Experimental Results The experimental results comparing single and double iterations of RL-based refinement are presented in Table 20.
@@
  • E.4 Proving the Effectiveness of Introducing Prediction Results in Refinement Paradigm
@@
    Appendix F Persona Probing
    F.1 Important Profiling Dimensions in Book Domain
@@
    F.2 Insights into User Group Characteristics
@@ G.2 Persona Evolution
    In this section, we display the evolution of the user’s persona and corresponding observations in each window over four persona update rounds with DeePer (Table  25, Table  26, Table  27, Table  28, Table  29).

### v2

@@ Abstract
To advance personalized applications such as recommendation systems and user behavior prediction, recent research increasingly adopts large language models (LLMs) for human-readable persona modeling. In dynamic real-world scenarios, effective persona modeling necessitates leveraging streaming behavior data to continually optimize user personas. However, existing methods—whether regenerating personas or incrementally extending them with new behaviors—often fail to achieve sustained improvements in persona quality or future behavior prediction accuracy. To address this, we propose DeePer, a novel approach for dynamic persona modeling that enables continual persona optimization. Specifically, we enhance the model’s direction-search capability through an iterative offline reinforcement learning framework, allowing it to automatically identify effective update directions and optimize personas using discrepancies between user behaviors and model predictions. Extensive experiments on dynamic persona modeling involving 4,800 users across 10 domains highlight DeePer ’s superior persona optimization capabilities, delivering an impressive 32.2% average reduction in user behavior prediction error over four update rounds—outperforming the best baseline by a remarkable 22.92%. [^1] Resources are available at https://github.com/sheep333c/DEEPER.git.
@@ 1 Introduction
Recent studies increasingly utilize Large Language Models (LLMs) OpenAI (2023a); Anthropic (2024); AI@Meta (2024) for human-readable and interpretable persona modeling, advancing personalized applications like recommendation and behavior prediction. However, most research focuses on generating personas from static historical data, which fail to capture dynamic behaviors and evolving preferences in real-world interactive scenarios Wang and Lim (2023); Zhou et al. (2024). This underscores the need for dynamic persona modeling—a pivotal yet underexplored approach that iteratively updates personas using streaming user behavior data to continually enhance their quality.
@@
To address the challenges, we propose DeePer (Directed Persona Refinement), a novel approach for LLM-based dynamic persona modeling. Specifically, we introduce a new paradigm, Persona Refinement (Figure 1), which uses discrepancies between user behaviors and model predictions as stronger update signals to expose deficiencies in personas. To identify effective update directions, we decompose the optimization objective into three direction search goals: Previous Preservation, Current Reflection, and Future Advancement, ensuring stability, adaptability, and task alignment. Based on these goals, we design reward functions for clear and measurable assessments of update directions by comparing predictive errors before and after updates. Finally, we propose an iterative offline reinforcement learning (RL) framework with two training stages, leveraging self-sampling and DPO fine-tuning to progressively enhance the model’s direction search and persona refinement capabilities, ultimately improving prediction accuracy.
@@ 2.1 Task Formulation
Consider a user $\mathcal{U}$ in domain $\mathcal{X}$. To capture temporal dynamics of user behaviors, we segment user’s online interactions into sequential, time-ordered windows $\mathbf{W}=\{\mathcal{W}_{t}\}_{t=0}^{\mathcal{T}}$. Each window $\mathcal{W}_{t}$ contains $N$ interactions, represented by an item list $\mathbf{I}_{t}=\{i_{t}^{j}\}_{j=1}^{N}$ and the corresponding user behaviors $\mathbf{O}_{t}=\{o_{t}^{j}\}_{j=1}^{N}$. As new data arrives at time $t$, the current window $\mathcal{W}_{t}$ captures interactions from the present period, while $\mathcal{W}_{t-1}$ reflects previous behaviors, and $\mathcal{W}_{t+1}$ outlines future interactions.
@@ 3.4 Iterative Training Framework
DeePer employs an iterative offline RL training framework (Figure 2): Iteration 1 fine-tunes the base model to refine initial personas (Model 1), while Iteration 2 further enhances it to refine pre-optimized personas (Model 2). Direct Preference Optimization (DPO)Rafailov et al. (2024) is used to seamlessly integrate rewards into preference pairs, enabling the model to identify better directions through explicit comparisons and supporting scalable iterative fine-tuning.
@@ 4.1 Experiment Setup
We use the frozen, powerful LLM, GPT-4o-miniOpenAI (2023a), to generate initial personas and predictions in a zero-shot setting for both training and evaluation, ensuring consistent initial persona quality and unbiased predictions. Additionally, it serves as the backbone for all baselines, offering a robust foundation for comparison.
@@ 5.1 What enables DeePer’s effectiveness
We first evaluate the necessity of direction search by comparing DeePer with frozen models (GPT-4o-mini and Llama-3.1-8B-Instruct (Base Model)) which refine personas directly. As shown in Figure 5(a), both baselines exhibit significant error increases after refinement, underscoring the critical role of direction search in effective optimization.
@@
Previous Window Prediction ($\varepsilon_{prev|\mathcal{S}_{old/new}}$) - Previous Preservation
@@
Figure 6: (a) Refinement performance across two offline RL iterations of DeePer. (b) Comparison of DeePer and fine-tuned IncUpdate (Inc-FT).
Iterative training enhances persona refinement.
@@ 5.2 Persona Probing
Refined personas from DeePer also enable in-depth, domain-specific exploration. In Book domain, we uncover group-level preferences by clustering final persona embeddings, identifying 5 user groups characterized by unique high-frequency adjectives (e.g., “romantic” and “practical”) (Figure 7(b)). We also extract domain-specific patterns by organizing high-frequency terms into 8 dimensions using GPT-4o OpenAI (2023b) (Table 3), highlighting critical factors for modeling Book domain users. These attempts show DeePer ’s potential to support strategic user insights exploration.
@@ 7 Conclusion
In this paper, we introduce DeePer, an effective approach to dynamic persona modeling that leverages iterative offline RL and discrepancy-based refinement to continuously enhance persona quality and predictive accuracy. Comprehensive experiments demonstrate DeePer ’s effectiveness across diverse domains in dynamic user modeling. We hope DeePer marks a significant advancement in personalized applications.
@@
Acknowledgement
Aili Chen, Yanghua Xiao et al. gratefully acknowledge the research funding provided by Alibaba Group. The authors also extend their sincere appreciation to Xintao Wang from Fudan University for his insightful feedback and valuable contributions throughout the research process.
@@ A.3 Persona Update
In this work, the formulation of persona update stage varies by corresponding paradigms. For all baselines, we use GPT-4o-mini as backbone, while for DeePer, we use the fine-tuned model via the iterative offline RL training framework based on Llama3.1-8b-Instruct. For each persona update method, we design corresponding update prompt template as follows.
@@
Amazon Review Data (2018)
- Movies and TV
Movies and TV Train/Eval 1000 837 E
MovieLens 20M Dataset
MovieLens Train/Eval 3000 1000 F
@@
Art Crafts and Sewing Eval - 86 G
@@
Sports and Outdoors Eval - 236 I
@@
Grocery and Gourmet Food Eval - 185 J
@@
A.4 User Details
@@
A.5 Training Data Construction
@@
In Iteration 2, Model 1 is first used to generate outputs for the 10,800 context data points from Iteration 1, completing the first persona update for each user. These results, in turn, are used to construct a second set of 10,800 context data points for the second persona refinement. These are then combined with the 10,800 context data points constructed in the first iteration, resulting in a total of 21,600 context data points for sampling in the second iteration. For each context, 15 candidate personas are randomly sampled using Model 1, with the same inference parameters as in Iteration 1. The boundaries for positive and negative reward sets are set to 0.5 and 0, with a margin of 0.8. A total of 28,612 new DPO preference pairs are generated. Additionally, 5,000 preference pairs with a margin greater than 0.8 are randomly selected from Iteration 1 to be included in the training set, mitigating the issue of catastrophic forgetting. This results in a total of 33,612 DPO preference pairs, with 10% randomly selected for the validation set, used to train Model 2.
@@
Appendix B Training Details
B.1 Hyperparameter Details
Our training pipeline is implemented based on LlamaFactory  Zheng et al. (2024) and the hyperparameters used in the training process are summarized in Table 12.
B.2 Loss Function Details
@@
The DPO loss optimizes the model by leveraging user preference signals to align persona refinements with higher rewards.
@@
Appendix C Dynamic Persona Modeling Task
@@
Grocery and Gourmet Food 1.19 1.20 1.16 1.14 1.13
@@
Art Crafts and Sewing 0.76 0.75 0.73 0.67 0.72
Automative 0.84 0.88 0.86 0.94 0.85
Book 1.00 1.03 0.98 0.97 0.93
@@
Grocery and Gourmet Food 1.19 1.19 1.16 1.18 1.14
@@
Movies and TV 1.12 1.12 1.11 1.10 1.08
Recipe 0.91 0.94 0.87 0.81 0.81
Sports and Outdoors 0.91 0.90 0.89 0.91 0.87
@@
Appendix D What Enables DeePer’s Effectiveness
@@
D.1 Proving Effectiveness of Direction search
@@
D.2 Proving the Effectiveness of Balanced Reward
@@
    D.3 Proving the Effectiveness of Iterative RL Training
    In this analysis, we aim to demonstrate the effectiveness of iterative offline RL training by comparing the performance of the model after one iteration of refinement versus two iterations. This comparison helps to understand whether additional refinement iterations contribute to improved model performance.
@@ D.3 Proving the Effectiveness of Iterative RL Training
    Experimental Results The experimental results comparing single and double iterations of refinement are presented in Table 20.
@@
  • D.4 Proving the Effectiveness of Introducing Prediction Results in Refinement Paradigm
@@
    Appendix E Persona Probing
    E.1 Important Profiling Dimensions in Book Domain
@@
    E.2 Insights into User Group Characteristics
@@
    Appendix F Additional Analysis
    F.1 Baselines Enhanced with Stronger Backbones and Task-Specific SFT
    To further validate the limitations of existing extension- and regeneration-based methods in dynamic persona modeling, we conduct additional experiments under two enhancement conditions:
    • (1) Stronger Backbone: All baselines are implemented with GPT-4o-mini as the underlying predictor and persona generator.
    • (2) Task-Specific SFT: We fine-tune baselines using task-specific supervision constructed via self-sampling and reject sampling, based on future prediction error comparisons.
    In both settings, we follow the same multi-round dynamic persona modeling task described in the main paper. Table F.1 summarizes the results after the first persona update.
    Table 24: Comparison of MAE before and after the first update across methods with strong backbones or task-specific SFT.
    Method Before Update After Update MAE Reduction
    DeePer 0.964 0.763 +20.9%
[…]
    HierMerge 0.964 0.974 -1.0%
    IncUpdate 0.964 0.924 +4.1%
    SlideRegen 0.964 0.996 -3.3%
    FullRegen 0.964 0.979 -1.6%
[…]
    HierMerge 0.964 0.985 -2.2%
    IncUpdate 0.964 0.963 +0.1%
    SlideRegen 0.964 1.025 -6.3%
    FullRegen 0.964 0.992 -2.9%
    As shown above, even when enhanced with stronger backbones or supervised fine-tuning, baseline methods fail to achieve meaningful optimization. In contrast, DeePer achieves a substantial MAE reduction. This provides further evidence that existing approaches suffer from a fundamental gap between persona updating and optimization, which DeePer effectively bridges through reward-guided direction refinement.
    F.2 Reusability and Scalability of DeePer Framework
    Although DeePer adopts a reinforcement learning-based training pipeline, which involves moderate computational overhead, we emphasize that its resulting dataset is reusable, model-agnostic, and adaptable to various downstream tasks. This significantly reduces the long-term costs of applying dynamic persona modeling in practice.
    1. Domain Generalization.
    The DeePer dataset is constructed using six training domains and generalizes effectively to unseen settings. As shown in the main results, DeePer achieves an average 36.4% MAE reduction across four unseen domains after four update rounds.
    2. Task Generalization.
    To evaluate task transferability, we construct a pair-wise prediction test on 600 users. After one refinement step, DeePer improves choice accuracy from 61% to 68%, reflecting an 11.4% increase in task-specific accuracy.
    Table 25: Pair-wise choice accuracy before and after refinement.
    Test Set Before After Accuracy Increase
    600 Users 0.61 0.68 +11.4%
    3. Model Generalization.
    We further apply the DeePer dataset to fine-tune a different backbone, Qwen2-7B-Instruct qwe (2024). The results demonstrate consistent gains: 13.1% MAE reduction in seen domains and 10.8% in unseen domains.
    Table 26: Performance of Qwen2-7B-Instruct fine-tuned on DeePer dataset.
    Domain Type Before MAE After MAE MAE Reduction
    Seen Domain 0.99 0.86 13.1%
    Unseen Domain 0.925 0.825 10.8%
    Conclusion.
    While DeePer’s offline RL training incurs moderate one-time cost, its dataset offers strong generalization across domains, tasks, and models—eliminating the need for repeated reward computation or full-model fine-tuning. This makes DeePer a practical and scalable solution for real-world dynamic persona modeling.
@@ G.2 Persona Evolution
    In this section, we display the evolution of the user’s persona and corresponding observations in each window over four persona update rounds with DeePer (Table  28, Table  29, Table  30, Table  31, Table  32).